Two masses 6.0 kg and 9.0 kg are hung from the two ends of a cord which passes over a smooth pulley so that the two masses hang freely. If initially the 9.0 kg mass is 4 meters from the floor, how long will it take this mass to reach the floor after the masses are released? 


1
Expert's answer
2021-07-09T08:31:45-0400
"a=\frac{9-6}{9+6}9.8=2.0\frac{m}{s^2}\\
0.5(2)t^2=4\\t=2\ s"